yomomalikebracobama:

Solve 2(x + 1) + 4 = 8. (1 point) a 1 b 6.5 c 4 d 3

yomomalikebracobama:

lets see

yomomalikebracobama:

this is easy im dum

yomomalikebracobama:

first distribute

yomomalikebracobama:

2x+2+4=8 so how i find x ?

QuestionCoveBot:

So you already distributed. Good. Secondly we have to add 2 and 4 to get 6. Then it looks like; 2x + 6 = 8

yomomalikebracobama:

where do we divide?

QuestionCoveBot:

The we subtract 6 from both sides. So subtract 6 from 8, and 6 from 6.

yomomalikebracobama:

ik we have to divide

QuestionCoveBot:

We don't divide.

yomomalikebracobama:

:(

QuestionCoveBot:

So 8-6 is 2. x=2

yomomalikebracobama:

2x+2

yomomalikebracobama:

den we divide

mxddi3:

Yes, divide. But it should say 2x=2

QuestionCoveBot:

Oh forgot that part. Uh yes, we divide. It looks like 2x = 2 so divide 2 and 2.

yomomalikebracobama:

2/2 is 1 or 2

QuestionCoveBot:

It's 1. So; x=1

yomomalikebracobama:

YES